# Detecting dilute axion stars constrained by fast radio bursts in the solar system via stimulated decay

### v1

@@ Abstract
Fast radio bursts (FRBs) can be explained by collapsing axion stars, imposing constraints on the axion parameter space and providing valuable guidance for experimental axion searches. In the traditional post-inflationary model, axion stars could constitute up to $75\%$ of the dark matter component, suggesting that some axion stars may exist within the solar system. Electromagnetic radiation with an angular frequency equal to half the axion mass can stimulate axion decay. Thus, directing a powerful radio beam at an axion star could trigger its stimulated decay, producing a detectable echo. Using this method, we find it is possible to test the existence of dilute axion stars with maximum masses ranging from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as constrained by FRBs, within the solar system. The resulting echo from axion stars constrained by FRBs could be detectable by terrestrial telescopes. Detecting such an echo would confirm the existence of axion stars, unravel the mystery of dark matter, and provide key evidence that some FRBs originate from collapsing axion stars. Furthermore, FRBs produced by axion star collapses could serve as standard candles, aiding in the resolution of the Hubble tension. If no echo is detected using this method, it would place constraints on the abundance of dark matter in the form of dilute axion stars with maximum masses in the range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$.
@@ I Introduction
Growing evidence from various observations and theoretical frameworks has established that dark matter constitutes a substantial portion of the Universe’s energy density. Nonetheless, its exact nature remains elusive. The QCD axion Weinberg:1977ma ; Wilczek:1977pj , arising from the Peccei-Quinn mechanism Peccei:1977ur ; Peccei:1977hh developed to resolve the strong CP problem, is a well-motivated candidate for dark matter. Additionally, string theory predicts a spectrum of axion-like particles (ALPs) across a wide mass range, a concept often referred to as the “axiverse” Arvanitaki:2009fg . In this article, we use the term “axions” to encompass both QCD axions and ALPs. Axions can be produced through several mechanisms, such as the misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah , the decay of string defects Gorghetto:2020qws , and the kinetic misalignment mechanism Co:2019jts . Due to their bosonic nature, axions can achieve extremely high phase space densities, leading to Bose-Einstein condensation (BEC) Sikivie:2009qn and the formation of gravitationally bound structures known as axion stars. For recent reviews, see Refs. Braaten:2019knj ; Visinelli:2021uve ; Zhang:2024bjo .
[…]
@@
Electromagnetic radiation with an angular frequency half the axion mass can stimulate the decay of axions, producing a detectable echo. This phenomenon offers a novel opportunity to search for axion dark matter. By emitting a powerful beam of microwave radiation into space and detecting the resulting echo, axion dark matter may be identified and studied Arza:2019nta ; Arza:2021nec ; Arza:2022dng ; Arza:2023rcs ; Gong:2023ilg . In our previous work, we proposed that the stimulated decay Kephart:1994uy ; Rosa:2017ury ; Caputo:2018vmy ; Di:2023nnb of a dilute axion star associated with planet 9 could be triggered by directing a powerful radio beam at the star, resulting in an echo detectable by terrestrial telescopes Di:2024snm . Such an observation would distinguish a dilute axion star from other planet 9 candidates, such as a PBH or a free-floating planet captured by the solar system Di:2024snm . This method offers a promising avenue for exploring the axion parameter space constrained by FRBs.
@@
This paper is organized as follows. In Sec. II, we briefly introduce the concept of dilute axion stars. In Sec. III, we review the constraints on the axion parameter space imposed by FRBs. Section IV discusses the abundance of dilute axion stars in the Universe. In Sec. V, we analyze the potential for detecting echoes from dilute axion stars, as constrained by FRBs, within the solar system by emitting powerful radio waves into these stars. Finally, Sec. VI presents our conclusions. Throughout this article, we use natural units, where $c=\hbar=1$.
@@ II Dilute Axion Stars
The QCD axion is a hypothetical pseudo-Nambu-Goldstone boson with spin-0, characterized by a small mass $m_{\phi}$, weak self-interactions, and extremely weak couplings to Standard Model particles. The invariance of the axion Lagrangian under the shift symmetry $\phi(x)\rightarrow\phi(x)+2\pi f_{a}$ requires the axion potential $V(\phi)$ to be periodic, such that $V(\phi)=V(\phi+2\pi f_{a})$, where $f_{a}$ is the axion decay constant, representing the energy scale at which the $U(1)_{\rm{PQ}}$ symmetry spontaneously breaks. Performing a series expansion of the axion potential around $\phi=0$, considering only the first two leading terms, the potential can be expressed as
@@
where $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$ represents the attractive self-interaction coupling constant. The value of $\kappa$ depends on the choice of axion potential. For the instanton potential, $V(\phi)=(m_{\phi}f_{a})^{2}\left[1-\cos\left({\phi}/{f_{a}}\right)\right]$, commonly used in axion phenomenology, $\kappa=1$. For the chiral potential DiVecchia:1980yfw ; GrillidiCortona:2015jxo , $\kappa\simeq 0.34$ GrillidiCortona:2015jxo ; Fujikura:2021omw . Considering the interaction between axions and photons, the general Lagrangian for the axion is given by
@@
where $F_{\mu\nu}$ is the electromagnetic field strength tensor, $\tilde{F}^{\mu\nu}$ is its dual, $\alpha$ is the fine-structure constant, and $K$ is a model-dependent constant of order one. Axions are not entirely stable and can decay into two photons via the interaction term ${\cal L}_{int}=\frac{\alpha K}{8\pi f_{a}}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$. This interaction has significant astrophysical implications and provides an experimental window for the search for these elusive particles. The decay rate of an axion at rest into two photons is
@@
For QCD axions, there exists a well-known relationship between the decay constant $f_{a}$ and the axion mass Weinberg:1977ma :
@@
where $m_{u}\simeq 2.2~{}\rm{MeV}$, $m_{d}\simeq 4.7~{}\rm{MeV}$, and $m_{\pi}\simeq 135~{}\rm{MeV}$ are the masses of the up quark, down quark, and pion, respectively, while $f_{\pi}\simeq 92~{}\rm{MeV}$ is the pion decay constant. This relationship is illustrated in Fig. 1. As bosons, axions can achieve exceptionally high phase space densities, potentially leading to the formation of BECs Sikivie:2009qn . These BECs can result in axion stars, which may manifest in both dilute and dense configurations Chavanis:2017loo ; Visinelli:2017ooc ; Eby:2019ntd . However, dense axion stars may have lifespans too short to significantly impact cosmology as astrophysical objects Braaten:2019knj ; Seidel:1991zh ; Hertzberg:2010yz ; Eby:2015hyx ; Wang:2020zur .
A stable dilute axion star is in equilibrium, where the attractive forces of self-gravity and self-interactions are balanced by the repulsive gradient energy. This equilibrium is maintained as long as the star’s density remains low, thereby minimizing the effects of self-interactions. However, if the mass of the dilute axion star exceeds a critical threshold, determined by the attractive self-interaction coupling constant, this equilibrium becomes disrupted. The maximum mass and corresponding minimum radius of a dilute axion star are given by Chavanis:2011zi ; Chavanis:2011zm
@@
where $M_{pl}$ is the Planck mass and $\lambda_{c}$ denotes the Compton wavelength of the axion. When the mass of a dilute axion star exceeds the critical mass given by Eq. (5), either due to merger events Mundim:2010hi ; Cotner:2016aaq ; Schwabe:2016rze ; Eby:2017xaw ; Hertzberg:2020dbk ; Du:2023jxh ; Maseizik:2024qly or through the accretion of axions from the surrounding environment Chen:2020cef ; Chan:2022bkz ; Dmitriev:2023ipv , self-interactions become significant, potentially leading to the collapse of the star Levkov:2016rkk ; Eby:2016cnq ; Fox:2023xgx . Substituting the attractive coupling constant of self-interaction, $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$, and the axion’s Compton wavelength, $\lambda_{c}=2\pi/m_{\phi}$, into Eq. (5), the critical mass and minimum radius are expressed as
@@ III Constraints by fast radio bursts
Since the axion is a pseudoscalar boson, the boson enhancement effect must be considered. The change in photon number density $n_{\lambda^{\prime}}$ within the dilute axion star is governed by the Boltzmann equation Kephart:1994uy :
@@
where $f_{\phi}$ and $f_{\lambda^{\prime}}$ represent the phase space densities of the axion and photon, respectively. The term $\cal M(\phi\rightarrow\gamma(\lambda^{\prime})\gamma(\lambda^{\prime}))$ refers to the matrix element corresponding to the interaction ${\cal L}_{\text{int}}=\frac{\alpha K}{8\pi f_{a}}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$. The phase space integration is performed using the standard Lorentz-invariant measure, considering the momenta of both the axion and the photons. Assuming the phase space distributions of axions and photons are approximately uniform and isotropic within the dilute axion star, the time evolution of photon number density becomes Kephart:1994uy ; Rosa:2017ury
$\frac{dn_{\gamma}}{dt}=\Gamma_{\phi}\left[2n_{\phi}\left(1+\frac{8\pi^{2}}{m_{\phi}^{3}v}n_{\gamma}\right)-\frac{16\pi^{2}}{3m_{\phi}^{3}}\left(v+\frac{3}{2}\right)n_{\gamma}^{2}\right],$
where $v\sim 1/(2R_{\rm{AS}}m_{\phi})$ represents the maximum axion velocity within the dilute axion star, as estimated using the Heisenberg uncertainty principle.
In Eq. (III), the initial terms within the square brackets represent both spontaneous and stimulated decay. The final term, scaled by a factor of $3/2$, corresponds to the production of “sterile” axions Kephart:1994uy , which can escape from the dilute axion star. Furthermore, photons produced through spontaneous and stimulated decay escape from the dilute axion star at a rate of $\Gamma_{e}=1/{R_{\rm{AS}}}$. By integrating Eq. (III) and accounting for the escaped photons, the coupled differential equations describing the evolution of the number of axions and photons within the dilute axion star are expressed as follows Di:2024tlz
@@
Figure 1: The parameter space of axions. The region above the orange solid line is excluded based on constraints from Planck 2018 combined with BAO measurements Caloni:2022uya . Constraints derived from white dwarf observations are represented by the green area Balkin:2022qer , with the latest findings detailed in Ref. Gomez-Banon:2024oux . The purple line corresponds to QCD axions, while the blue region represents the parameter space constrained by FRBs, where the maximum mass of dilute axion stars ranges from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$.
Interestingly, radio signals from the stimulated decay of collapsing axion stars may already have been detected. FRBs are bright, transient radio emissions lasting only milliseconds Lorimer:2007 ; Keane:2012yh ; Thornton:2013iua , with a frequency range spanning approximately 400 MHz to 8 GHz Petroff:2019tty and total energy outputs typically between $10^{38}$ and $10^{40}$ erg Thornton:2013iua ; Spitler:2014fla . The photon frequency and total energy emitted by a collapsing axion star, for certain parameter ranges, align with those observed in FRBs Di:2024tlz . For constraints on the axion parameter space derived from FRBs, see Ref. Di:2024tlz and Fig. 1. The maximum axion star mass within this constrained parameter space ranges from approximately $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as shown in Fig. 1.

### v4

@@ Abstract
Fast radio bursts (FRBs) can be explained by collapsing axion stars, imposing constraints on the axion parameter space and providing valuable guidance for experimental axion searches. In the traditional post-inflationary model, axion stars could constitute up to $75\%$ of the dark matter component, suggesting that some axion stars may exist within the Solar System. Photons with energy half the axion mass can stimulate axion decay. Thus, directing a powerful radio beam at an axion star could trigger its stimulated decay, producing a detectable echo. Using this method, we find it is possible to test the existence of dilute axion stars with maximum masses ranging from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as constrained by FRBs, within the Solar System. The resulting echo from axion stars constrained by FRBs could be detectable by terrestrial telescopes. Detecting such an echo would confirm the existence of axion stars, unravel the mystery of dark matter, and provide key evidence that some FRBs originate from collapsing axion stars. Furthermore, FRBs produced by axion star collapses could serve as standard candles, aiding in the resolution of the Hubble tension. If no echo is detected using this method, it would place constraints on the abundance of dark matter in the form of dilute axion stars with maximum masses in the range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$.
@@ I Introduction
Growing evidence from various observations and theoretical frameworks has established that dark matter constitutes a substantial portion of the Universe’s energy density. Nonetheless, its exact nature remains elusive. The QCD axion Weinberg:1977ma ; Wilczek:1977pj , arising from the Peccei-Quinn mechanism Peccei:1977ur ; Peccei:1977hh developed to resolve the strong-$CP$ problem, is a well-motivated candidate for dark matter. Additionally, string theory predicts a spectrum of axion-like particles (ALPs) across a wide mass range, a concept often referred to as the “axiverse” Arvanitaki:2009fg . In this article, we use the term “axions” to encompass both QCD axions and ALPs. Axions can be produced through several mechanisms, such as the misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah ; Co:2019jts , thermal production Turner:1986tb ; Salvio:2013iaa , the decay of string Davis:1986xc ; Gorghetto:2020qws , and primordial black hole (PBH) evaporation Schiavone:2021imu ; Li:2022xqh ; Mazde:2022sdx . Because axions are bosons, they can attain very high phase space densities, which may result in Bose-Einstein condensation (BEC) Sikivie:2009qn and the formation of gravitationally bound objects known as axion stars Braaten:2019knj ; Visinelli:2021uve ; Gorghetto:2024vnp ; Chang:2024fol ; Zhang:2024bjo .
The unusual orbits of trans-Neptunian objects (TNOs) Brown:2004yy ; Trujillo ; Batygin:2016zsa and the gravitational anomalies detected by the Optical Gravitational Lensing Experiment (OGLE) Mroz:2017mvf remain unexplained. One potential explanation is the existence of PBHs Niikura:2019kqi or axion stars Sugiyama:2021xqg with masses in the range of $M\sim 0.5-20M_{\oplus}$, which could account for the OGLE observations. To address the TNO orbital anomalies, the planet 9 hypothesis proposes an object with a mass of $5-15M_{\oplus}$ (approximately $1.5-4.5\times 10^{-5}M_{\odot}$) located $300-1000$ AU from the Sun Batygin-PR . This hypothetical planet 9 could be explained as a dilute axion star Di:2023xaw that was captured by the Solar System. Additionally, collapsing axion stars with specific parameters may emit millisecond-long radio bursts with peak luminosities of approximately $10^{42}$ erg/s, which are consistent with the characteristics of observed non-repeating fast radio bursts (FRBs) Di:2023nnb . These collapsing axion stars could serve as novel standard candles Di:2024tlz for constraining the Hubble constant, $H_{0}$, due to their strong, intrinsic luminosity being fixed and dependent solely on the axion mass and decay constant. Other studies have also investigated the relationship between axion stars and FRBs Iwazaki:2014wka ; Tkachev:2014dpa ; Raby:2016deh ; Buckley:2020fmh . If some FRBs are indeed produced by collapsing axion stars, this would impose constraints on the axion parameter space Di:2024tlz , providing valuable guidance for experimental searches for axions. Thus, exploring the axion parameter space in relation to FRBs is crucial, as it offers insight into the origin of FRBs and the nature of dark matter.
@@
Photons with energy half the axion mass can trigger the stimulated decay Kephart:1994uy ; Rosa:2017ury ; Caputo:2018vmy ; Di:2023nnb ; Dev:2023ijb of axions, producing a echo. This phenomenon offers a novel opportunity to search for axion dark matter. By emitting a radio beam with high power into space and detecting the resulting echo, axion dark matter may be identified and studied Arza:2019nta ; Arza:2021nec ; Arza:2022dng ; Arza:2023rcs ; Gong:2023ilg . In our previous work Di:2024snm , we proposed that the stimulated decay of a dilute axion star associated with planet 9 could be triggered by directing a powerful radio beam at the star, resulting in an echo could be detected by current radio telescopes. Such an observation would distinguish an axion star from a PBH Scholtz:2019csj or other planet 9 candidates. This method offers a promising avenue for exploring the axion parameter space constrained by FRBs.
@@
This article is organized as follows. In Sec. II, we briefly introduce the concept of dilute axion stars. In Sec. III, we review the constraints on the axion parameter space imposed by FRBs. Section IV discusses the abundance of dilute axion stars in the Universe. In Sec. V, we analyze the potential for detecting echoes from dilute axion stars, as constrained by FRBs, within the Solar System by emitting powerful radio beams into these stars. Finally, Sec. VI presents our conclusions. In this article, we adopt natural units, setting $c=\hbar=1$.
@@ II Dilute Axion Stars
The QCD axion is a hypothetical pseudo-Nambu-Goldstone boson with spin-0, characterized by a small mass $m_{\phi}$, weak self-interactions, and extremely weak couplings to Standard Model particles. The invariance of the axion Lagrangian under the shift symmetry $\phi(x)\rightarrow\phi(x)+2\pi f_{a}$ requires the axion potential $V(\phi)$ to be periodic, such that $V(\phi)=V(\phi+2\pi f_{a})$, where $f_{a}$ represents the $U(1)$ symmetry breaking scale, also known as the axion decay constant. Performing a series expansion of the axion potential around $\phi=0$, and considering only the first two leading terms, the potential can be expressed as
@@
where $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$ is the coupling constant representing the attractive self-interaction of axions. The value of $\kappa$ depends on the choice of axion potential. For the instanton potential, $V(\phi)=(m_{\phi}f_{a})^{2}\left[1-\cos\left({\phi}/{f_{a}}\right)\right]$, commonly used in axion phenomenology, $\kappa=1$. For the chiral potential DiVecchia:1980yfw ; GrillidiCortona:2015jxo , $\kappa\simeq 0.34$ GrillidiCortona:2015jxo ; Fujikura:2021omw . Considering the interaction between axions and photons, the general Lagrangian for the axion is given by
@@
where $\alpha$ is the fine-structure constant representing the electromagnetic field coupling strength, $K$ is a dimensionless quantity of order one that depends on the axion model, and $F_{\mu\nu}$ and $\tilde{F}^{\mu\nu}$ represent the electromagnetic field tensor and its dual, respectively. Due to their coupling with electromagnetic fields, axions can spontaneously decay into photons. This interaction has significant astrophysical implications and provides an experimental avenue for searching for these elusive particles. The spontaneous decay rate of an axion at rest into two photons is given by
@@
The extremely low decay rate makes it challenging to detect axions through photon signals from their spontaneous decay. For QCD axions, there exists a well-known relationship between the decay constant $f_{a}$ and the axion mass Weinberg:1977ma :
@@
where $m_{u}\simeq 2.2~{}\rm{MeV}$, $m_{d}\simeq 4.7~{}\rm{MeV}$, and $m_{\pi}\simeq 135~{}\rm{MeV}$ are the masses of the up quark, down quark, and pion, respectively, while $f_{\pi}\simeq 92~{}\rm{MeV}$ is the pion decay constant. This relationship is illustrated in Fig. 1. As bosons, axions can achieve exceptionally high phase space densities, potentially leading to the formation of BECs Sikivie:2009qn and the creation of axion stars, which are divided into dilute and dense branches Schiappacasse:2017ham ; Chavanis:2017loo ; Visinelli:2017ooc ; Eby:2019ntd . Dilute axion stars are stable against perturbations, while dense axion stars may have lifetimes much shorter than any cosmological timescale Braaten:2019knj ; Seidel:1991zh ; Hertzberg:2010yz ; Eby:2015hyx ; Wang:2020zur .
For a stable dilute axion star, equilibrium is maintained by the balance between attractive self-gravity and self-interactions, and the repulsive pressure from the Heisenberg uncertainty principle. This equilibrium persists as long as the star’s density remains low, minimizing the influence of self-interactions. However, once the star’s mass exceeds a critical threshold, determined by the strength of the attractive self-interaction coupling constant, the equilibrium is disrupted. The maximum mass $M_{\rm{max}}$ and corresponding minimum radius $R_{\rm{min}}$ of a stable axion star are given by Chavanis:2011zi ; Chavanis:2011zm
@@
where $\lambda_{c}$ is the Compton wavelength of the axion and $M_{pl}$ is the Planck mass. When the mass of an axion star exceeds the critical mass given by Eq. (5), either due to merger events Mundim:2010hi ; Cotner:2016aaq ; Schwabe:2016rze ; Eby:2017xaw ; Hertzberg:2020dbk ; Du:2023jxh ; Maseizik:2024qly or through the accretion of axions from the surrounding environment Chen:2020cef ; Chan:2022bkz ; Dmitriev:2023ipv , self-interactions become significant, potentially leading to the collapse of the star Levkov:2016rkk ; Chavanis:2016dab ; Eby:2016cnq ; Fox:2023xgx . Substituting the attractive coupling constant of self-interaction, $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$, and the axion’s Compton wavelength, $\lambda_{c}=2\pi/m_{\phi}$, into Eq. (5), the critical mass and minimum radius are expressed as
@@ III Constraints by fast radio bursts
Despite the extremely low spontaneous decay rate, axion decay rates can be enhanced in the presence of ambient photons, as photons with energy equivalent to half of the axion mass can stimulate decays of axions into two photons. The change in photon number density $n_{\lambda^{\prime}}$ within a dilute axion star, resulting from axion decay and inverse decay, is governed by the Boltzmann equation Kephart:1994uy :
@@
where $\cal M(\phi\rightarrow\gamma(\lambda^{\prime})\gamma(\lambda^{\prime}))$ refers to the matrix element corresponding to the coupling term between axions and photons, $f_{\phi}$ is the phase space density of axions, and $f_{\lambda^{\prime}}$ is the phase space density of photons with helicity $\lambda^{\prime}$. After performing the phase space integration for the above equation and considering the two possible helicities of photons, the time evolution of the total photon number density becomes Kephart:1994uy ; Rosa:2017ury
$\frac{dn_{\gamma}}{dt}=2\Gamma_{\phi}n_{\phi}+\frac{16\pi^{2}}{m_{\phi}^{3}v}\Gamma_{\phi}n_{\phi}n_{\gamma}-\frac{16\pi^{2}}{3m_{\phi}^{3}}\left(v+\frac{3}{2}\right)\Gamma_{\phi}n_{\gamma}^{2},$
where $v$ is the maximum velocity of an axion within the axion star, which is roughly $1/(2R_{\rm{AS}}m_{\phi})$ as estimated using the uncertainty principle.
Figure 1: Constraints on the axion parameter space. The region above the orange line is excluded based on Planck 2018 and BAO measurements Caloni:2022uya . A recent study Springmann:2024ret places a model-independent bound on QCD axions from supernova 1987A, which is stronger than the bounds from Planck and BAO. Constraints from white dwarf observations are represented by the green region Balkin:2022qer , with recent updates provided in Ref. Gomez-Banon:2024oux . The purple line indicates constraints for QCD axions, while the blue region corresponds to the parameter space constrained by FRBs, where the maximum mass of dilute axion stars ranges from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$.
In Eq. (III), the first term represents spontaneous axion decay, while the second term accounts for the stimulated decay of axions in the presence of ambient photons. The third term corresponds to the inverse decay process, in which photons are converted back into axions. The component with a factor of $3/2$ corresponds to the generation of “sterile” axions with high velocity that escape from the axion star Kephart:1994uy . Additionally, photons produced through spontaneous axion decay and stimulated decay in the presence of ambient photons can also escape from the axion star, with an escape rate of $\Gamma_{e}=1/{R_{\rm{AS}}}$. Integrating Eq. (III) and considering the photons that escape from the axion star, the coupled differential equations describing the evolution of the number of axions and photons within the axion star are given by Di:2024tlz
@@
Interestingly, radio signals from the stimulated decay of collapsing axion stars may already have been detected. FRBs are bright, transient radio emissions lasting only milliseconds Lorimer:2007 ; Keane:2012yh ; Thornton:2013iua , with a frequency range spanning approximately 400 MHz to 8 GHz Petroff:2019tty and total energy outputs typically between $10^{38}$ and $10^{40}$ erg Thornton:2013iua ; Spitler:2014fla . The photon frequency and total energy emitted by a collapsing axion star, for certain parameter ranges, align with those observed in FRBs Di:2024tlz . For constraints on the axion parameter space derived from FRBs, see Ref. Di:2024tlz and Fig. 1. The constraints on the axion parameter space provide valuable guidance for experimental searches for axions. The maximum axion star mass within this constrained parameter space, assuming $\kappa=1$, ranges from approximately $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as shown in Fig. 1.